OpenStudy (anonymous):

Niki is building a kite. She has a rectangular piece of cloth that is 11 inches wide and 15 inches long. She needs to place a stick on the diagonal of the rectangle to make her kite. How long of a stick does she need? Round to the nearest tenth of an inch. 10.2 inches 26.0 inches 18.6 inches 4.0 inches

OpenStudy (anonymous):

draw a rectangle and split it diagonally, into two triangles, now use a^2+b^2=c^2 to find c^2 (which is one of the multiple answers)

OpenStudy (anonymous):

can you just give me the answer?please

OpenStudy (anonymous):

11^2 + 15^2 = c^2 121 + 225 = 346^2 (unsquare 346) 346/18.6= 18.6 The answer is 18.6